Given a 0.584 liter sample of gas collected over water at 22 degrees Celcius and 636 torr, find the volume occupied by the dry gas at STP
1
Expert's answer
2016-06-01T14:31:03-0400
pV=nRT
R=62.363577 L·Torr K−1·mol−1
K = °C + 273
Even being under different conditions, the amount of gas remains the same:
n=pV/RT=(636·0.584)/(62.363·(22+273))=0.02 moles
STP: temperature of 273.15 K and an absolute pressure of exactly 100 kPa (1 bar, 14.504 psi, 0.98692 atm).
100 kPa = 750 torr
V=nRT/p
V=(0.02·62.363·273.15)/750=0.45 liters